# Highly accurate simulations of asymmetric black-hole scattering and cross validation of effective-one-body models

### v1

@@ III.1.1 EOB dynamics
The dissipative effects enter the dynamics through the radiation-reaction force $\mathcal{F}$, which are computed as a summation of PN GW modes (augmented with gravitational self-force information van de Meent et al. (2023)) in a factorised form Damour et al. (2009); Damour and Nagar (2007); Pan et al. (2011); Damour and Nagar (2008); Pompili et al. (2023). For arbitrary orientations of the spins, both the spins and the orbital plane precess around the total angular momentum of the system $\bm{J}$. However, in this work we restrict to spins aligned with the orbital angular momentum of the systems. For such non-precessing spins the equations of motion read Buonanno et al. (2006); Khalil et al. (2023)
@@
where the evolution of the radial momentum employs the transformation to tortoise-coordinates, $p_{r^{*}}=p_{r}\xi(r)$, with $\xi(r)=dr/dr^{*}$ (see Refs. Taracchini et al. (2014); Pan et al. (2014) for details).
@@ III.1.2 SEOBNRv5 Hamiltonian
where the two terms in the Hamiltonian include the even-in-spin and the odd-in-spin contributions, respectively. The term $H_{\rm even}$ contains the non-spinning and spin-spin contributions through the potentials $A^{\text{align}}(r)$, $B_{np}^{\text{align}}(r)$ and $Q^{\text{align}}(r)$, with no S4 corrections needed since the Kerr Hamiltonian reproduces all even-in-spin leading PN orders for BBHs Vines and Steinhoff (2018). The potentials include 4PN spin-spin information,
@@
where the $A_{\text{noS}}$ term is Taylor-expanded up to 5PN in $u\equiv M/r$ with the replacement of the coefficient $u^{6}$ (except for the log term), by the parameter $a_{6}$ which is calibrated to NR simulations. The explicit expressions for the rest of the potentials can be found in Ref. Khalil et al. (2023) and in Appendix A of Ref. Pompili et al. (2023).
The odd-in-spin term, $H_{\rm odd}$, contains the gyromagnetic factors, $g_{a_{+}}$ and $g_{a_{-}}$, expanded up to 3.5PN Khalil et al. (2023), the function $G_{a^{3}}^{\text{align}}$ which includes $S^{3}$ corrections (see Appendix A of Ref. Pompili et al. (2023) for the explicit expressions) and a calibration term at 4.5PN,
@@ III.1.3 Modifications for scattering
One advantage of the evolved model is that they can also simulate dynamical capture scenarios. This allows them to determine the critical parameters where the system is captured as well as the maximum value of the scattering angle on the separatrix. This is in contrast to the closed-form models where there is no maximum value of the scattering angle, only the divergence of the model close to the separatrix.
@@ III.2 Closed-form EOB
$p_{r}^{2}$ $=\frac{1}{(1+B_{\rm np}^{\rm Kerr})}\bigg{[}\frac{1}{A}\left(E_{\rm eff}-\frac{ML(g_{a_{+}}a_{+}+g_{a_{-}}\delta a_{-})}{r^{3}+a_{+}^{2}(r+2M)}\right)^{2}$
@@
this implicitly defines the effective potential $w_{\rm\texttt{SEOB-PM}{}}$. The function $A$ is
@@
and $A^{\rm Kerr}$, $B_{\rm npa}^{\rm Kerr}$, $B_{\rm np}^{\rm Kerr}$ are all parameters of the Kerr metric. Perturbative PM deformations are incorporated into $\Delta A$ and $\Delta g_{a_{\pm}}$ in a similar fashion to Eq. (23). Full details are given in Ref. Buonanno et al. (2024b).
In both closed-form models, the PM deformation coefficients, $w^{(m)}$ and $\Delta A$, $\Delta g_{a_{\pm}}$, are uniquely determined by the perturbative PM coefficients of the scattering angle. More precisely, they are polynomial functions of each other; the analytic complexity of the closed-form potentials is thus equivalent to the PM scattering angle. The $n$PM scattering-angle coefficients have a simple form. First, they are $\lfloor\tfrac{n-1}{2}\rfloor$SF exact where $n$SF denotes an expansion in the symmetric mass ratio $\nu$ to the $n$th order (i.e. 1PM and 2PM are 0SF exact, 3PM and 4PM are 1SF exact and so on). Second, non-trivial analytic dependence appears only as functions of the effective energy $E_{\rm eff}$. At leading 1PM and 2PM orders all dependence is rational, and at higher $n$PM orders multiple polylogarithms of weight $(n-2)$ Driesse et al. (2024b) appear (at 5PM order, the spinless 2SF part is still unknown). Additionally at the 4PM order, complete elliptic integrals appear Bern et al. (2021).
While the closed-form models seem conservative by construction, one may include dissipative effects by considering different choices of $\theta$ in Eq. (22). For a conservative model, one would use the conservative analytic PM angle. However, one may also match the potentials against dissipative angles. With dissipation, the effects of recoil generally imply that the scattering angles of each BH differ, and one may then consider different choices for $\theta$. These include the relative scattering angle $\theta_{\rm rel}$ Bini et al. (2021); Driesse et al. (2024b) and the individual scattering angles $\theta_{i}$ of each BH.
The relative scattering angle commonly used in PN and PM literature is defined as the angle between incoming and outgoing relative momenta, $\bm{p}_{\rm in}$ and $\bm{p}_{\rm out}$, in the initial and final CoM frames. That is, importantly, the outgoing spatial momentum $\bm{p}_{\rm out}$ is defined in the final CoM frame — which differs from the initial CoM frame due to radiation of four momentum of the two black holes. Its computation thus involves the recoil vector $\bm{P}_{\rm recoil}$ describing the back-reaction of the CoM frame — for more detail, see e.g. Refs. Dlapa et al. (2023a); Driesse et al. (2024b). The relative angle is then simply the angle between the incoming and outgoing momenta:
@@
The relative scattering angle is symmetric in the two BHs, but it is also different from the averaged angle (14):
@@
Thus, for each potential, we have four subclasses of models depending on which angle they were matched against: a conservative potential matched from the conservative angle, a dissipative potential matched from the (dissipative) relative scattering angle and two asymmetric models, one for each $\theta_{i}$, matched from the individual scattering angles. Further, the models are labelled by the precision to which they reproduce $\theta$. Namely, for each model (conservative, relative and asymmetrical) we have a $n$PM version with $n\leq 5$. Note, however, that dissipative effects start only at 3PM such that all angles agree until then. Further, recoil effects start only at 4PM (unequal masses) and 5PM (unequal spins) such that the asymmetric models only distinguish themselves there.
The 5PM scattering angle is known only to the 1SF accuracy with its 2SF ($\nu^{2}$) piece still undetermined. As such, the 5PM models considered in this work —namely SEOB-5PM(1SF)— are matched only at 1SF accuracy. Technically, and for consistency, we include $G^{5}$ 2SF terms to 3rd PN order which are determined from the potential itself via the 4PM results (generally $n$PM fully determines $(n-1)$PN observables through the potential). These 2SF 3PN terms must be included in the 5PM angle in order that the matched potential is free of poles in its PN limit.
The asymmetric models have not previously been considered and are introduced in this work to capture recoil effects on the scattering angle. Again, there are two asymmetric models built to reproduce the scattering angle $\theta_{i}$ of each individual BH. The potentials of each model are determined by inserting analytic $n$PM expressions for the individual scattering angle of the $i$th BH $\theta_{i}$ in Eq. (22) and enforcing the validity of this equation to the $n$th PM order (for an $n$PM order model). In particular, using these models we may compute a prediction for the angle difference $\theta_{2}-\theta_{1}$. Namely, we compute $\theta_{1}$ using the first asymmetrical model and $\theta_{2}$ using the second one and calculate their difference.
Dissipative models based on the relative scattering angle accurate to the fourth (physical) PM order were constructed in Refs. Rettegno et al. (2023); Buonanno et al. (2024b) (with partial 5PM spin deformations). A model including conservative 5PM 1SF deformations (based on the PM results of Ref. Driesse et al. (2024a)) was considered in Ref. Swain et al. (2025). However, the recent 5PM dissipative results of Ref. Driesse et al. (2024b) have not yet been tested in EOB models.
@@ IV Scattering-angle comparisons
In this section we compare the scattering angles calculated from SpEC with those obtained using the EOB models described in Sec. III, as well as the publically available evolution model TEOBResumS-Dalí [^2] In particular, we compare to v.1.1.0 available at https://bitbucket.org/teobresums/teobresums/src/v1.1.0-Dali/.. One major difference between the evolution models is that SEOBNRv5 only includes quasicircular information in the radiation-reaction terms, while TEOBResumS-Dalí includes analytical non-circular corrections to the azimuthal radiation-reaction force Nagar et al. (2025); Chiaramello and Nagar (2020).
@@
For each simulation we provide the initial conditions, SpEC scattering angle (with errors), and the relative scattering angle calculated using each EOB model in Appendix B. This data, as well as the exact initial parameters of the simulations, are also available in the ancillary data file.
@@ IV.1 Equal-mass non-spinning black holes
The left panel of Fig. 6 shows the comparison of the scattering angle for the sequence of simulations at constant energy $\Gamma$. Most of the EOB models show agreement with the NR results to $<3\%$ right up to capture, which matches previous comparisons in the literature Damour et al. (2014); Rettegno et al. (2023); Swain et al. (2025). The exception to this is TEOBResumS-Dalí which diverges from the NR results as we decrease $\ell$ and predicts that the separatrix is at a larger $\ell$. In contrast, the SEOBNRv5 model stays faithful to the NR results up to the NR capture point, but it predicts that the systems stay unbound until a lower $\ell$ value. Due to the increased accuracy of the SpEC results we can look at the behaviour of the EOB models when approaching the weak field. The inset in the lower panel shows that the EOB models are all tending towards the NR values as we increase $\ell$ apart from TEOBResumS-Dalí which appears to slightly overshoot.
@@
In Fig. 7, we explore the effects of higher order PM contributions by comparing the SEOB-4PM and SEOB-5PM(1SF) models. Evidently, the agreement of the SEOB-4PM with NR is best. This could be explained by the missing 5PM(2SF) contributions in the SEOB-5PM(1SF) model because, naively, all SF orders are similarly important for the equal-mass phase space considered here. In addition to the SEOB-PM models, we have also plotted the perturbative PM angles (i.e. without any resummation) at 4PM and 5PM precision including, either, only 0SF contributions (4PM(0SF) and 5PM(0SF)) or 0SF and 1SF contributions (4PM and 5PM(1SF)). Here, it seems that the 1SF contributions pull the (geodesic) 0SF angles away from the NR result worsening the agreement. However, when one considers exact 0SF results, the angle overshoots and from this perspective the 1SF contributions should decrease the angle slightly (see Fig. 4 of Ref. Driesse et al. (2024b)). We also note the interesting feature, that at the perturbative level, 1SF corrections are very small compared with 0SF. However, the resummation models are very sensitive even to these very small variations.
Figure 7: Scattering angle results for equal mass systems with $\Gamma=1.055$ and comparisons to higher order PM models. In the legend, $n$PM has all terms up to $G^{n}$ whereas $n$PM($m$SF) has all terms up to $G^{n-1}$ plus contributions to $G^{n}$ up to $\nu^{m}$. The vertical dotted line shows the first confirmed capture of the SpEC simulations.
@@ IV.2 Equal-mass aligned-spin black holes
The closed-form models show a very similar behaviour to the NR results and are generally within a few percent of the SpEC values. The evolution models both predict a capture for $\chi_{+}\lesssim-0.1$, which starts at significantly lower spin than the SpEC capture range ${\chi_{+}\lesssim-0.26}$. SEOBNRv5 agrees with the NR results in the non-spinning limit, whereas TEOBResumS-Dalí tends towards the SpEC values when increasing $\chi_{+}$. This behaviour is consistent with the results of Ref. Hopper et al. (2023) where an earlier version of TEOBResumS also showed disagreement with NR results exceeding 10% for aligned-spin systems, as well as, predicting premature captures.
Figure 10: Comparison of EOB scattering angle results with those calculated by SpEC for unequal-mass non-spinning systems. The top panels show the relative scattering angle for constant $\gamma\!=\!1.02$, $\ell\!=\!4.8$ (left) and $\gamma\!=\!1.09158$, $b\!=\!10.5M$ (right). The bottom panels show the relative difference in the relative scattering angle as compared to the NR values. The vertical dotted lines on the right plot show the first confirmed capture.
The right panel of Fig. 8 shows the scattering angle for the equal-magnitude, anti-parallel spin simulations. The upper plot shows that increasing the spin magnitudes has little effect on the scattering angle. The lower panel shows that the agreement between the EOB models does not vary greatly when altering the spin magnitudes. The closed-form models agree with the NR results to $\lesssim 5\%$, whereas the SEOBNRv5HM and TEOBResumS-Dalí models have relative differences to the NR of $\sim 9\%$ and $\sim 55\%$ respectively.
Fig. 9 shows the individual scattering angles $\theta_{i}$ for each BH for the anti-parallel simulations. From the top panel, we see that the scattering angle across the range of spins is almost identical, only varying by fractions of a percent. Nevertheless, the difference in the scattering angles between the two BHs is apparent. The lower plot shows that the difference between the two angles grows as the magnitude of the spins increases the asymmetric emission in the system. Although previous NR results for unequal aligned spins exist Rettegno et al. (2023); Albanesi et al. (2025b), this work is the first to directly measure the BHs’ scattering angle difference, which, being $<0.1\%$ of the total angle, requires the high accuracy of the SpEC results. The small magnitude of this difference is unsurprising, as this effect emerges only at 5PM order and is therefore extremely subdominant Buonanno et al. (2024b).

### v2

@@ III.1.1 EOB dynamics
The dissipative effects enter the dynamics through the radiation-reaction force $\mathcal{F}$, which are computed as a summation of PN GW modes (augmented with gravitational self-force information van de Meent et al. (2023)) in a factorized form Damour et al. (2009); Damour and Nagar (2007); Pan et al. (2011); Damour and Nagar (2008); Pompili et al. (2023). For arbitrary orientations of the spins, both the spins and the orbital plane precess around the total angular momentum of the system $\bm{J}$. However, in this work we restrict to spins aligned with the orbital angular momentum of the systems. For such non-precessing spins the equations of motion read Buonanno et al. (2006); Khalil et al. (2023)
@@
where the evolution of the radial momentum employs the transformation to tortoise-coordinates, $p_{r_{*}}=p_{r}\xi(r)$, with $\xi(r)=dr/dr^{*}$ (see Refs. Taracchini et al. (2014); Pan et al. (2014) for details).
@@ III.1.2 SEOBNRv5 Hamiltonian
where $a_{\pm}=a_{1}\pm a_{2}$. The two terms in the Hamiltonian include the even-in-spin and the odd-in-spin contributions, respectively. The term $H_{\rm even}$ contains the non-spinning and spin-spin contributions through the potentials $A^{\text{align}}(r)$, $B_{np}^{\text{align}}(r)$ and $Q^{\text{align}}(p_{r_{*}},r)$, with no $S^{4}$ corrections needed since the Kerr Hamiltonian reproduces all even-in-spin leading PN orders for BBHs Vines and Steinhoff (2018). The potentials include 4PN spin-spin information,
@@
where the $A$, $B$, and $D$ potentials are functions of $r$ and the $Q$ potentials depend on $r$ and $p_{r_{*}}$. The $A_{\text{noS}}$ term is Taylor-expanded up to 5PN in $u\equiv M/r$ with the replacement of the coefficient $u^{6}$ (except for the log term), by the parameter $a_{6}$ which is calibrated to NR simulations. The explicit expressions for the potentials can be found in Ref. Khalil et al. (2023) and in Appendix A of Ref. Pompili et al. (2023).
The odd-in-spin term, $H_{\rm odd}$, contains the gyromagnetic factors, $g_{a_{+}}(p_{\phi},r)$ and $g_{a_{-}}(p_{\phi},r)$, expanded up to 3.5PN Khalil et al. (2023), the function $G_{a^{3}}^{\text{align}}(p_{\phi},r)$ which includes $S^{3}$ corrections (see Appendix A of Ref. Pompili et al. (2023) for the explicit expressions) and a calibration term at 4.5PN,
@@ III.1.3 Modifications for scattering
One advantage of the evolved models is that they can also simulate dynamical capture scenarios. This allows them to determine the critical parameters where the system is captured as well as the maximum value of the scattering angle on the separatrix. This is in contrast to the closed-form models where there is no maximum value of the scattering angle, only the divergence of the model close to the separatrix.
@@ III.2 Closed-form EOB
$p_{r}^{2}$ $=\frac{1}{(1+B_{\rm np}^{\rm Kerr})}\bigg[\frac{1}{A}\left(E_{\rm eff}-\frac{ML(g_{a_{+}}a_{+}+g_{a_{-}}\delta\>a_{-})}{r^{3}+a_{+}^{2}(r+2M)}\right)^{2}$
@@
this implicitly defines the effective potential $w_{\rm\texttt{SEOB-PM}{}}$. The function $A$ and the gyro-gravitomagnetic factors $g_{a_{\pm}}$ are
@@
and $A^{\rm Kerr}(r)$, $B_{\rm npa}^{\rm Kerr}(r)$, and $B_{\rm np}^{\rm Kerr}(r)$ are all parameters of the Kerr metric. Perturbative PM deformations are incorporated into $A(E_{\rm eff},\chi_{\pm},r)$ and $g_{a_{\pm}}(E_{\rm eff},\chi_{\pm},r)$ through $\Delta A(E_{\rm eff},\chi_{\pm},r)$ and $\Delta g_{a_{\pm}}(E_{\rm eff},\chi_{\pm},r)$ in a similar fashion to Eq. (23). Full details are given in Ref. Buonanno et al. (2024b).
In both closed-form models, the PM deformation coefficients, $w^{(m)}$ and $\Delta A$, $\Delta g_{a_{\pm}}$, are uniquely determined by the perturbative PM coefficients of the scattering angle. More precisely, they are polynomial functions of each other; the analytic complexity of the closed-form potentials is thus equivalent to the PM scattering angle. The nnPM scattering-angle coefficients have a simple form. First, they are $\lfloor\tfrac{n-1}{2}\rfloorSF$ exact where nnSF denotes an expansion in the symmetric mass ratio $\nu$ to the $n$th order (i.e. 1PM and 2PM are 0SF exact, 3PM and 4PM are 1SF exact and so on). Second, non-trivial analytic dependence appears only as functions of the effective energy $E_{\rm eff}$. At leading 1PM and 2PM orders all dependence is rational, and at higher nnPM orders multiple polylogarithms of weight (n2)(n-2) Driesse et al. (2025) appear (at 5PM order, the spinless 2SF part is still unknown). Additionally at the 4PM order, complete elliptic integrals appear Bern et al. (2021).
While the closed-form models seem conservative by construction, one may include dissipative effects by considering different choices of $\theta$ in Eq. (22). For a conservative model, one would use the conservative analytic PM angle. However, one may also match the potentials against dissipative angles. With dissipation, the effects of recoil generally imply that the scattering angles of each BH differ, and one may then consider different choices for $\theta$. These include the relative scattering angle $\theta_{\rm rel}$ Bini et al. (2021); Driesse et al. (2025) and the individual scattering angles $\theta_{i}$ of each BH.
The relative scattering angle commonly used in PN and PM literature is defined as the angle between incoming and outgoing relative momenta, $\bm{p}_{\rm in}$ and $\bm{p}_{\rm out}$, in the initial and final CoM frames. That is, importantly, the outgoing spatial momentum $\bm{p}_{\rm out}$ is defined in the final CoM frame — which differs from the initial CoM frame due to radiation of four momentum of the two black holes. Its computation thus involves the recoil vector $\bm{P}_{\rm recoil}$ describing the back-reaction of the CoM frame — for more detail, see e.g. Refs. Dlapa et al. (2023a); Driesse et al. (2025). The relative angle is then simply the angle between the incoming and outgoing momenta:
@@
For a covariant version of this formula we refer to Eq. (26) of Ref. Jakobsen et al. (2024). The relative scattering angle is symmetric in the two BHs, but it is also different from the averaged angle (14):
@@
Thus, for each potential, we have four subclasses of models depending on which angle they were matched against: a conservative potential matched from the conservative angle, a dissipative potential matched from the (dissipative) relative scattering angle and two asymmetric models, one for each $\theta_{i}$, matched from the individual scattering angles. Further, the models are labeled by the precision to which they reproduce $\theta$. Namely, for each model (conservative, relative and asymmetrical) we have a nnPM version with $n\leq 5$. Note, however, that dissipative effects start only at 3PM such that all angles agree until then. Further, recoil effects start only at 4PM (unequal masses) and 5PM (unequal spins) such that the asymmetric models only distinguish themselves there.
The 5PM scattering angle is known only to the 1SF accuracy with its 2SF ($\nu^{2}$) piece still undetermined. As such, the 5PM models considered in this work —namely SEOB-5PM(1SF)— are matched only at 1SF accuracy. Technically, and for consistency, we include $G^{5}$ 2SF terms to 3rd PN order which are determined from the potential itself via the 4PM results (generally nnPM fully determines (n1)(n-1)PN observables through the potential). These 2SF 3PN terms must be included in the 5PM angle in order that the matched potential is free of poles in its PN limit. More details on the SEOB-5PM(1SF) model are given in Appendix B.
The asymmetric models have not previously been considered and are introduced in this work to capture recoil effects on the scattering angle. Again, there are two asymmetric models built to reproduce the scattering angle $\theta_{i}$ of each individual BH. The potentials of each model are determined by inserting analytic nnPM expressions for the individual scattering angle of the $i$th BH $\theta_{i}$ in Eq. (22) and enforcing the validity of this equation to the $n$th PM order (for an nnPM order model). In particular, using these models we may compute a prediction for the angle difference $\theta_{2}-\theta_{1}$. Namely, we compute $\theta_{1}$ using the first asymmetrical model and $\theta_{2}$ using the second one and calculate their difference.
Dissipative models based on the relative scattering angle accurate to the fourth (physical) PM order were constructed in Refs. Rettegno et al. (2023); Buonanno et al. (2024b) (with partial 5PM spin deformations). A model including conservative 5PM 1SF deformations (based on the PM results of Ref. Driesse et al. (2024)) was considered in Ref. Swain et al. (2025). However, the recent 5PM dissipative results of Ref. Driesse et al. (2025) have not yet been tested in EOB models.
@@ IV Scattering-angle comparisons
In this section we compare the scattering angles calculated from SpEC with those obtained using the EOB models described in Sec. III, as well as the publically available evolution model TEOBResumS-Dalí [^2] In particular, we compare to v.1.1.0 available at https://bitbucket.org/teobresums/teobresums/src/v1.1.0-Dali/.. One major difference between the evolution models is that SEOBNRv5 only includes quasicircular information in the radiation-reaction terms, while TEOBResumS-Dalí includes analytical non-circular corrections to the azimuthal radiation-reaction force Nagar et al. (2025); Chiaramello and Nagar (2020). Generally, unless otherwise stated, we use the closed-form models based on the relative angle which include all dissipative effects encoded in the relative angle.
@@
For each simulation we provide the initial conditions, SpEC scattering angle (with errors), and the relative scattering angle calculated using each EOB model in Appendix C. This data, as well as the exact initial parameters of the simulations, are also available in the ancillary data file.
@@ IV.1 Equal-mass non-spinning black holes
The left panel of Fig. 6 shows the comparison of the scattering angle for the sequence of simulations at constant energy $\Gamma$. Most of the EOB models show agreement with the NR results to $3\%$ right up to capture, which matches previous comparisons in the literature Damour et al. (2014); Rettegno et al. (2023); Swain et al. (2025). The exception to this is TEOBResumS-Dalí which diverges from the NR results as we decrease $\ell$ and predicts that the separatrix is at a larger $\ell$. In contrast, the SEOBNRv5 model stays faithful to the NR results up to the NR capture point, but it predicts that the systems stay unbound until a lower $\ell$ value. Due to the increased accuracy of the SpEC results we can look at the behavior of the EOB models when approaching the weak field. The inset in the lower panel shows that the EOB models are all tending towards the NR values as we increase $\ell$ apart from TEOBResumS-Dalí which appears to slightly overshoot.
@@
In Fig. 7, we explore the effects of higher order PM contributions by comparing the SEOB-4PM and SEOB-5PM(1SF) models. Evidently, the agreement of the SEOB-4PM with NR is best. This could be explained by the missing 5PM(2SF) contributions in the SEOB-5PM(1SF) model because, naively, all SF orders are similarly important for the equal-mass phase space considered here. In addition to the SEOB-PM models, we have also plotted the perturbative PM angles (i.e. without any resummation) at 4PM and 5PM precision including, either, only 0SF contributions (4PM(0SF) and 5PM(0SF)) or 0SF and 1SF contributions (4PM and 5PM(1SF)). Here, it seems that the 1SF contributions pull the (geodesic) 0SF angles away from the NR result worsening the agreement. However, when one considers exact 0SF results, the angle overshoots and from this perspective the 1SF contributions should decrease the angle slightly (see Fig. 4 of Ref. Driesse et al. (2025)). We also note the interesting feature, that at the perturbative level, 1SF corrections are very small compared with 0SF. However, the resummation models are very sensitive even to these very small variations.
Figure 7: Scattering angle results for equal mass systems with $\Gamma=1.02264$ and comparisons to higher order PM models. In the legend, nnPM has all terms up to $G^{n}$ whereas nnPM(mmSF) has all terms up to $G^{n-1}$ plus contributions to $G^{n}$ up to $\nu^{m}$. The vertical dotted line shows the first confirmed capture of the SpEC simulations.
@@ IV.2 Equal-mass aligned-spin black holes
The closed-form models show a very similar behavior to the NR results and are generally within a few percent of the SpEC values. The evolution models both predict a capture for $\chi_{+}\lesssim-0.1$, which starts at significantly lower spin than the SpEC capture range $\chi_{+}\lesssim-0.26$}. SEOBNRv5 agrees with the NR results in the non-spinning limit, whereas TEOBResumS-Dalí tends towards the SpEC values when increasing $\chi_{+}$. This behavior is consistent with the results of Ref. Hopper et al. (2023) where an earlier version of TEOBResumS also showed disagreement with NR results exceeding 10% for aligned-spin systems, as well as, predicting premature captures.
Figure 10: Comparison of EOB scattering angle results with those calculated by SpEC for unequal-mass non-spinning systems. The top panels show the relative scattering angle for constant $\gamma\!=\!1.02$, $\ell\!=\!4.8$ (left) and $\gamma\!=\!1.0916$, b=10.5Mb\!=\!10.5M (right). The bottom panels show the relative difference in the relative scattering angle as compared to the NR values. The vertical dotted lines on the right plot show the first confirmed capture.
The right panel of Fig. 8 shows the scattering angle for the equal-magnitude, anti-parallel spin simulations. The upper plot shows that increasing the spin magnitudes has little effect on the scattering angle, as previously observed in previous studies of similar systems Rettegno et al. (2023). The lower panel shows that the agreement between the EOB models does not vary greatly when altering the spin magnitudes. The closed-form models agree with the NR results to $\lesssim 5\%$, whereas the SEOBNRv5HM and TEOBResumS-Dalí models have relative differences to the NR of $\sim 9\%$ and $\sim 55\%$ respectively.
Fig. 9 shows the individual scattering angles $\theta_{i}$ for each BH for the equal magnitude, anti-parallel simulations. From the top panel, we see that the scattering angle across the range of spins is almost identical, only varying by fractions of a percent. Nevertheless, the difference in the scattering angles between the two BHs is apparent. The lower plot shows that the difference between the two angles grows as the magnitude of the spins increases the asymmetric emission in the system. Although previous NR results for unequal aligned spins exist Rettegno et al. (2023); Albanesi et al. (2025b), this work is the first to directly measure the BHs’ scattering angle difference, which, being $0.1\%$ of the total angle, requires the high accuracy of the SpEC results. The small magnitude of this difference is unsurprising, as this effect emerges only at 5PM order and is therefore extremely subdominant Buonanno et al. (2024b).